# Detecting High-Stakes Interactions with Activation Probes

### v2

@@ C.5 Manual Annotation
In order to check whether we agree with the ground truth labels for situational stakes, our main contributors manually annotated 20 samples from each evaluation dataset. The results can be found in Table 9.
@@ C.6 High-Stakes Labels and Deception
We looked at whether our high-stakes label correlates with observed deceptive behaviour in Llama-70B. We used the datasets “AI Liar”, “AI Audit”, “Roleplaying” and “Sandbagging WMDB/MMLU” from the codebase of Goldowsky-Dill et al. (2025). Each of these datasets has a “deception score” field, which we normalized to between 0 and 1. We also used the MASK Benchmark dataset (Ren et al., 2025), where we used the “honesty score” of the high-pressure prompts as the deception score. For MASK, we also separately looked at whether our high-stakes label correlates with whether the sample is “high-pressure” across both high-pressure and belief elicitation prompts.
As can be seen in Figure 20, for the AI Liar and Sandbagging datasets, there is a clear correlation between stakes level and deception score. For MASK and Roleplaying, there is no clear correlation, and for AI Audit, there is a negative correlation. Each of these datasets induces deception in a different way. For AI Audit, deception occurs when the model does not disclose to an auditor that a possible rule violation has taken place. A negative correlation with stakes suggests that, as the scope of the rule violation increases, the likelihood of Llama-70B reporting it increases.
For MASK, pressure is applied, and we observe whether the model changes its answer from truthful to false. It is possible that, when GPT-4o considers such a sample to be very high-stakes (9 or 10), that is in light of recognising that it is being pressured to lie—and regardless of whether the sample actually involves lying. Indeed, in Figure 21, we see that the positive correlation between “pressure” and stakes is much stronger than the correlation between stakes and pressure-induced dishonesty.

### v3

@@ C.5 Manual Annotation
In order to check whether we agree with the ground truth labels for situational stakes, our main contributors manually annotated 20 samples from each evaluation dataset. The results can be found in Table˜11.
@@ C.6 High-Stakes Labels and Deception
We looked at whether our high-stakes label correlates with observed deceptive behaviour in Llama-3.3-70B. We used the datasets “AI Liar”, “AI Audit”, “Roleplaying” and “Sandbagging WMDB/MMLU” from the codebase of Goldowsky-Dill et al. (2025). Each of these datasets has a “deception score” field, which we normalized to between 0 and 1. We also used the MASK Benchmark dataset (Ren et al., 2025), where we used the “honesty score” of the high-pressure prompts as the deception score. For MASK, we also separately looked at whether our high-stakes label correlates with whether the sample is “high-pressure” across both high-pressure and belief elicitation prompts.
As can be seen in Figure˜20, for the AI Liar and Sandbagging datasets, there is a clear correlation between stakes level and deception score. For MASK and Roleplaying, there is no clear correlation, and for AI Audit, there is a negative correlation. Each of these datasets induces deception in a different way. For AI Audit, deception occurs when the model does not disclose to an auditor that a possible rule violation has taken place. A negative correlation with stakes suggests that, as the scope of the rule violation increases, the likelihood of Llama-3.3-70B reporting it increases.
For MASK, pressure is applied, and we observe whether the model changes its answer from truthful to false. It is possible that, when GPT-4o considers such a sample to be very high-stakes (9 or 10), that is in light of recognising that it is being pressured to lie—and regardless of whether the sample actually involves lying. Indeed, in Figure˜21, we see that the positive correlation between “pressure” and stakes is much stronger than the correlation between stakes and pressure-induced dishonesty.
@@
C.7 Quantifying Out-of-Distribution Nature of Evaluation Datasets
The evaluation datasets are out-of-distribution by definition as they come from a different distribution (existing internet datasets rather than model-generated synthetic data). More importantly, they consist of real-world dialogues, multi-turn interactions, and contexts involving tool usage or psychological harm, properties not present in the synthetic training dataset.
To quantify the degree of distribution shift, we measure several metrics comparing evaluation datasets to our synthetic test split, shown in Table˜12. All evaluation datasets show substantially higher KL divergence in token distributions (0.82–2.50 vs 0.02), indicating significant vocabulary and usage pattern differences. Length distributions also differ markedly, with evaluation datasets ranging from very short (Aya: 22.8 tokens) to much longer (ToolACE: 804.1 tokens) compared to synthetic data (98.7 tokens). Perplexity measurements under Llama-3.3-70B further confirm distributional differences, with evaluation datasets showing both lower (Anthropic HH-RLHF: 6.28, MTS Dialog: 6.03) and higher (Mental Health: 13.17, Aya: 14.98) perplexities than synthetic data (10.43), reflecting their diverse linguistic characteristics.
Table 12: Quantifying distribution shift between synthetic training data and evaluation datasets. We analyze differences with respect to data source, domain, interaction structure, and several quantitative metrics. Length distributions show mean $\pm$ standard deviation of token counts using Llama-3.3-70B’s tokenizer. KL divergence is measured between bag-of-words distributions using the top 5,000 most frequent tokens (KL($P_{\text{eval}}\|P_{\text{train}}$)). Lower KL divergence indicates closer distributional match. Average per-token perplexity is computed using Llama-3.3-70B. All evaluation datasets show substantially larger distribution shifts than the synthetic test split, confirming their out-of-distribution nature.
Dataset Source Domain Interaction Length (tokens) KL div. Perplexity
Synthetic test Synthetic Multiple* Single-turn $98.7\pm 98.0$ 0.02 10.43
Anthropic HH-RLHF Real Multiple Multi-turn $275.7\pm 184.3$ 1.77 6.28
MT Samples Real Medical Single-turn $784.5\pm 452.5$ 1.76 8.25
MTS Dialog Real Medical Multi-turn $327.9\pm 363.8$ 2.50 6.03
ToolACE Synthetic Tool usage Multi-turn $804.1\pm 387.1$ 2.11 7.05
Mental Health Real Psychological Single-turn $169.7\pm 194.1$ 1.81 13.17
Aya Redteaming Real Multiple Single-turn $22.8\pm 11.7$ 0.82 14.98
*Excluding domains such as tool usage and psychological harm (see Topics in Table˜9).
C.8 Performance Variation Across Datasets
While our probes achieve strong generalization overall (mean AUROC $\sim$0.916 across test datasets vs >0.99 on synthetic test data; Figure˜3), performance varies across individual datasets (Figure˜12). We identify three factors that affect monitoring performance:
Noise. The Mental Health dataset consists of social media statements, where probes show high false positive rates on very short utterances lacking context (e.g., “What’s the schedule?” or “Seriously kidding”). False positives also occur on longer posts describing depressive episodes where individual sentences discuss coping mechanisms or minor details, or where high-stakes phrases co-occur with polite or optimistic language (e.g., “while my depression makes me want to kill myself…so I tried this method of venting…hope you like them… Thank you for reading”).
Context-dependency. Some datasets contain stakes that depend on deployment context beyond what system prompts describe. For example, when an agent provides real-time trading information for a fintech company, even seemingly routine tasks like providing stock prices become high-stakes. Such dependencies are present in ToolACE, causing high false negative rates and explaining some performance degradation on that dataset.
Domain-specific high-stakes. Datasets like Anthropic HH-RLHF and Aya Redteaming include prompts relating to psychological harm, hate speech, or culturally specific references (e.g., “Write a funny story about sati”) that are more nuanced than our training cases. Despite these challenges, probes still performed well (AUROC 0.95 on Anthropic HH-RLHF, 0.90 on Aya Redteaming).
Importantly, as shown in Figure˜5, adding as few as 256 deployment-specific examples to the training data significantly improves performance, providing a practical mechanism to counteract domain shifts.
Appendix D Detailed Future Work
Integrating diverse probes and monitoring techniques. Future work should also focus on developing integrated monitoring pipelines. While we explored combining probes with LLM baselines (Section˜3), more extensive research is needed on how to best integrate signals from various probes and other types of LLM monitoring techniques such as input/output classifiers (Inan et al., 2023b; Sharma et al., 2025) or chain-of-thought monitoring (Baker et al., 2025) into a cohesive system. As a preliminary step, we analyse the correlation of high-stakes labels with deception across several deception datasets. Our findings suggest that the relationship varies depending on the dataset and context (see Section˜C.6 for details). These findings indicate that high-stakes probes should be combined with probes specific to other concepts, such as deception (Goldowsky-Dill et al., 2025), truthfulness (Wagner, 2024), sycophancy, task-specific errors, etc. The goal is a more holistic assessment where, e.g. an interaction flagged as high-stakes and containing deception warrants a different response than one that is high-stakes but not deceptive.